OpenStudy (anonymous):

how would i find the vertex,axis of symmetry, y and x intercepts, max/min value, range of 2x^2+5x-3?

OpenStudy (anonymous):

Did we go over this earlier?

OpenStudy (anonymous):

yep haha, but I'm preparing for a test this time.

OpenStudy (nenadmatematika):

people don't trust you rickjbr hahha :D

OpenStudy (anonymous):

lol guess not

OpenStudy (anonymous):

The vertex is at \[(\frac{-b}{2a},f(\frac{-b}{2a}))\]

OpenStudy (anonymous):

In this form, that is, ax^2+bx+c

OpenStudy (anonymous):

The axis of symmetry is at the x-coordinate of the vertex, the max or min is at the y coordinate of the vertex

OpenStudy (anonymous):

To find the y-intercepts, set x=0. To find the x-intercepts, set y=0.

OpenStudy (anonymous):

ax^2+bx+c if a < 0, the y-coordinate of the vertex is a maximum because the parabola opens down. if a>0, the y-coordinate of the vertex is a minimum because the parabola opens up.

OpenStudy (anonymous):

You should try to find everything yourself. Others will review your work.

OpenStudy (anonymous):

i understand most of the ideology behind this, but can your further explain how i get the vertex part

OpenStudy (anonymous):

Tell me what a, b, and c are

OpenStudy (nenadmatematika):

ouchhhhhh :D

OpenStudy (anonymous):

Have to start from the beginning, don't we?

OpenStudy (anonymous):

a=2, b=5, c=-3

OpenStudy (anonymous):

Right, so what is: \[\frac{-b}{2a}\]

OpenStudy (anonymous):

-5/4

OpenStudy (anonymous):

good, that's the x-coordinate of the vertex.

OpenStudy (anonymous):

how do we find the y-coordinate?

OpenStudy (anonymous):

plug -5/4 into x?

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

so what'd you come up with?

OpenStudy (anonymous):

sec pls haha

OpenStudy (anonymous):

50/8-25/4-3?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

how would you do it?

OpenStudy (anonymous):

\[2(\frac{-5}{4})^2+5(\frac{-5}{4})-3\]

OpenStudy (anonymous):

how would you simplify that then?

OpenStudy (anonymous):

\[2(\frac{-5}{4})^2-\frac{25}{4}-3=2(\frac{25}{16})-\frac{25}{4}-3\]

OpenStudy (anonymous):

\[=\frac{50}{16}-\frac{25}{4}-3\]

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

-50/16-3?

OpenStudy (anonymous):

That's equivalent to the correct answer, but you can simplify it more

OpenStudy (anonymous):

-98/16?

OpenStudy (anonymous):

Keep going.

OpenStudy (anonymous):

What is -98/16 simplified?

OpenStudy (anonymous):

49/8

OpenStudy (anonymous):

-49/8

OpenStudy (anonymous):

my bad, but i get that.

OpenStudy (anonymous):

So what's the vertex

OpenStudy (anonymous):

(-5/4, -49/8)

OpenStudy (anonymous):

good, what's the axis of symmetry?

OpenStudy (anonymous):

-5/4

OpenStudy (anonymous):

x=-5/4

OpenStudy (anonymous):

Does this function have a maximum value or minimum value?

OpenStudy (anonymous):

mi

OpenStudy (anonymous):

n

OpenStudy (anonymous):

how do you know?

OpenStudy (anonymous):

2 is positive

OpenStudy (anonymous):

right, so what is the minimum value

OpenStudy (anonymous):

same as vertex

OpenStudy (anonymous):

the minimum value is not a point

OpenStudy (anonymous):

? how so

OpenStudy (anonymous):

the minimum value is -49/8

OpenStudy (anonymous):

ohhhh, min y value. ok ok

OpenStudy (anonymous):

now, what's the range of this function?

OpenStudy (anonymous):

(-infinity, -49/8]U[-49/8, infinity)

OpenStudy (anonymous):

does this graph ever reach negative infinity?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

then what's the range?

OpenStudy (anonymous):

i gave you the range though...?

OpenStudy (anonymous):

but it's wrong

OpenStudy (anonymous):

Hint: the function has a minimum value

OpenStudy (anonymous):

both would be positive (-49/8, infinity)?

OpenStudy (anonymous):

Is the vertex included in the range?

OpenStudy (anonymous):

not sure. vertex the zero of the equation?

OpenStudy (anonymous):

Is the minimum value included in the range, instead?

OpenStudy (anonymous):

The function is defined at the minimum... Range: [-49/8, infinity)

OpenStudy (anonymous):

got it.

OpenStudy (anonymous):

ok, what are the y-intercepts

OpenStudy (anonymous):

what is the y-intercept***, there is only 1

OpenStudy (anonymous):

3

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

right, x=0 at the y-intercept. The y-intercept is then (0,-3) Don't forget it is a point

OpenStudy (anonymous):

What are the x-intercepts?

OpenStudy (anonymous):

x-intercepts im not sure about.

OpenStudy (anonymous):

OK, in this particular example, we can factor it. This isn't always the case, so you might need the quadratic formula

OpenStudy (anonymous):

2x^2+5x-3 2x^2+6x-x-3 Can you factor from here?

OpenStudy (anonymous):

This is the grouping method

OpenStudy (anonymous):

unfamiliar with the grouping method. but x(2x+5)

OpenStudy (anonymous):

Do you know the quadratic formula?

OpenStudy (anonymous):

Any way you get the x-intercepts is fine, be it factoring or using the quadratic formula

OpenStudy (anonymous):

Can you factor 2x^2+5x-3 without grouping?

OpenStudy (anonymous):

(2x )(x )

OpenStudy (anonymous):

sorry, got distracted.

OpenStudy (anonymous):

so, can you get the x-intercepts?

OpenStudy (anonymous):

trial and error is fine, this can be factored

OpenStudy (anonymous):

lol. rickjbr, something came up i extremely appreciate your help. thanks for going through the steps with mua.

OpenStudy (anonymous):

ill come back to this when i have the chance. But thanks you again.